OpenStudy (anonymous):

Solve for the variable. 5 - x = 16

OpenStudy (anonymous):

5x^2-4= 16 x^2-9x+20=0 9/10 = C/48 3x^2+x-1=0 and 9 - 3x < 4 (-x+1)

OpenStudy (agreene):

5-x=16 subtract 5 from both sides -x=11 multiply both by -1 x=-11

OpenStudy (anonymous):

5x + 2y = 16 ------------- (1) 3x + 4y = 18 ------------- (2) eq(1) X 2 - eq(2) we get, 10x + 4y - 3x - 4y = 32 -18 7x = 14 x =2 putting the value of x in eq(1) 10 + 2y =16 2y = 6 y =3 therefore, x = 2 and y = 3 Edited 6 months ago

OpenStudy (karatechopper):

x=-11

OpenStudy (anonymous):

no

OpenStudy (karatechopper):

then what? thts ur question

OpenStudy (anonymous):

we have to substitute x and y in two ways and subtract

OpenStudy (anonymous):

the 5

OpenStudy (karatechopper):

u should have said tht in ur post

OpenStudy (anonymous):

u to

OpenStudy (anonymous):

i am is write

OpenStudy (anonymous):

u know

OpenStudy (anonymous):

u r moron

OpenStudy (anonymous):

karatechopper

OpenStudy (anonymous):

your name is funny

OpenStudy (anonymous):

idiot

OpenStudy (anonymous):

dont fight

OpenStudy (karatechopper):

im not fighting geez... just stating the facts

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we are friends

OpenStudy (anonymous):

i like you

OpenStudy (karatechopper):

yay!

OpenStudy (anonymous):

idiot

OpenStudy (karatechopper):

:(

OpenStudy (anonymous):

give my medal i gave u

OpenStudy (anonymous):

giv

OpenStudy (karatechopper):

i will give if u stop calling me mean names

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

give

OpenStudy (anonymous):

thank you

OpenStudy (karatechopper):

welcome:)

OpenStudy (anonymous):

sorry

OpenStudy (karatechopper):

it ok:)

OpenStudy (anonymous):

for typing the dirty words

OpenStudy (karatechopper):

it fine

OpenStudy (karatechopper):

lol nice

OpenStudy (anonymous):

thanks